12. [-/1 Points] DETAILS SCALCET8 A.E.044. Find the limit. ..om :11 +1] Need Help? Read It Talk to a Tutor Submit Answer
Answer #1

Solution: lim z n too [CH]$1] ial we get using Reimman suum f(x) = 5(x²+) a = 0, b=1 = 5 (2+1) dx = 5 562+x]. 5(4+1) = 25 ans